OpenStudy (anonymous):

q if p, q, r, s are non-zero numbers p, q are the roots of x^2 + rx + s r, s are the roots of x^2 + px + q what is the value of p + q + r + s?

OpenStudy (anonymous):

@dan815 @amistre64 @hartnn

hartnn (hartnn):

Sum of roots for \(\Large ax^2+bx+c\) will be \(\Large -\dfrac{b}{a}\)

hartnn (hartnn):

so sum of roots for x^2 + rx + s = ... ?

OpenStudy (anonymous):

-r

OpenStudy (anonymous):

= p + q

hartnn (hartnn):

yes

OpenStudy (anonymous):

also r + s = -p

hartnn (hartnn):

thats correct

OpenStudy (anonymous):

p + q + r + s = -(r + p)?

OpenStudy (anonymous):

i'm not able to proceed do i have to plug in the values of p or q and r or s into the equation?

hartnn (hartnn):

i would say -(r+p) is the final answer...is a numeric answer expected?

OpenStudy (anonymous):

yes, a numeral.

hartnn (hartnn):

p + q + r + s = -r+r+s = s similarly, p + q + r + s = -p+p+q = q which means s=q hmm

OpenStudy (dan815):

p, q are the roots of x^2 + rx + s r, s are the roots of x^2 + px + q (x-p)(x-q)=x^2+rx+s x^2+(-q-p)x+qp=x^2+rx+s -q-p=r qp=s (x-r)(x-s)=x^2+px+q similarily -r-s=p sr=q ------------------4 system of equations -q-p=r qp=s -r-s=p sr=q

OpenStudy (dan815):

4 unknowns, linearly independant so solved

hartnn (hartnn):

with s=q p=r=1 and -(p+r) =-2

OpenStudy (anonymous):

p + q + r + s = s 2 + q + s = s q = -2 = s

hartnn (hartnn):

yes, we allready knew p + q + r + s = -(r+p) = -2

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^